Some compounds have interatomic bonds that are partially ionic and partially covalent. Calculate the percentage ionic character of the following compounds: HF, HCl and HBr. Use this information to explain why HF has a higher boiling temperature than HCl (19.4 vs. -85oC), even though HF has a lower molecular weight.
Answer #1

The below image provides the solution.
Answer #2

Step 1: Calculate % Ionic Character

Use Pauling's formula:

%Ionic=(1e0.25(ΔEN)2)×100

Electronegativity (EN): H (2.20), F (3.98), Cl (3.16), Br (2.96).

  • HF:

    ΔEN=3.982.20=1.78%Ionic=(1e0.25(1.78)2)×100=43.2%

  • HCl:

    ΔEN=3.162.20=0.96%Ionic=(1e0.25(0.96)2)×100=19.7%

  • HBr:

    ΔEN=2.962.20=0.76%Ionic=(1e0.25(0.76)2)×100=12.4%


Step 2: Why HF Boils Higher Than HCl

  • HF has 43% ionic character → forms strong hydrogen bonds (H-F is very polar).

  • HCl only has 20% ionic character → weaker dipole forces (no H-bonding).

  • Even though HF is lighter, H-bonds require more energy to break, so it boils at 19.4°C vs. HCl at -85°C.


Key Formula Used:

%Ionic=(1e0.25(ΔEN)2)×100

Result: HF > HCl > HBr in ionic character and boiling point.


answered by: anonymous